OpenStudy (anonymous):

Solve (2x+1)(x-1)=14

OpenStudy (anonymous):

2x*x + 2x*-1 + 1*x + 1*-1 = 14 2x^2 -x -15 = 0

OpenStudy (anonymous):

(2x+1)(x-1)=14 2x^2-x-1=14 -14 -14 2x^2-x-15=0 Apply quadratic equation if you need to find the zeros of the equation.

OpenStudy (solomonzelman):

x=3 or x= -5/2